### v2

@@ 4.1 Selection of Data Sources
We select Stack Overflow as a primary data source for our study. It is the largest programming Q&A site for programming topics containing over 24 million questions and 35 million answers SOUsageData . Thus, Stack Overflow could be a potential source for deep learning (DL) bugs. Developers often submit their encountered problems on Stack Overflow, when building their deep learning applications. According to a recent work zhao2021state , Stack Overflow contains at least 30 topics and 80K posts related to deep learning issues, which makes it a potential source for our data. We also select the benchmark dataset of Moravati et al. defects4ml (Defects4ML in Fig. 2) containing 100 DL bugs as our second source of data.
@@
We combine all four filtration criteria above, construct a SQL query, and then execute the query against the Stack Exchange Data Explorer. The SQL query is available in our replication package replicationPackage . After this filtration step, we obtained a total of 279 posts for Keras, 1,433 posts for Tensorflow, and 1,700 posts for PyTorch. To ensure the validity of our chosen filtration criteria and that the posts are related to DL bugs, we conducted a follow-up manual analysis on a representative sample of 75 bugs. We select the sample for the manual analysis using a 90% confidence interval and a 10% margin of error. We analyze the entire Stack Overflow discussion and code changes in GitHub for each of these bugs, and we found that 96% of them discuss deep learning bugs, confirming the validity of our filtration criteria (Steps 1-2, Fig. 2).
@@ 4.4 Qualitative Analysis
Before performing any qualitative analysis, we check if each selected post accurately represents the type of bug it is assigned to. To validate the bug categorization, we manually analyze each selected Stack Overflow post and review the issue description, stack trace, observed behaviour, expected behaviour, and accepted answer. This ensures that the tag-based selection of posts from Stack Overflow does not affect our results’ validity. We perform this manual verification since the Stack Overflow posts might lack an independently verified category label. In contrast, the GitHub issues in the benchmark dataset by Morovati et al. defects4ml already contain a validated category label.
Once we confirm the category of each post or issue, we follow a two-step approach to reproduce the deep learning bugs from Stack Overflow posts and Github issues. First, we gather complementary information about each bug, such as the dataset, code snippet, library versions, and the framework used. Second, we attempt to reproduce the bug using the code snippet and supporting data (e.g., dataset information, environment configurations, hyperparameters, and training logs). The first author and one independent collaborator conducted the bug reproduction process. We first reproduced 10 bugs from our dataset and achieved a Cohen Kappa of 54.5%. Then, we had two meetings to identify the main reasons for our disagreements and resolved them. In the next round, we reproduced 10 more bugs and achieved a Cohen Kappa of 89.1%, which is considered an almost perfect agreement mchugh2012interrater . After achieving an almost perfect agreement, the first author reproduced the remaining bugs (i.e., 128) while the independent collaborator checked the reproduction, achieving an average Cohen Kappa of 85.4%. If the first author fails to reproduce the bug within 60 minutes, the independent collaborator also attempts to reproduce the bug. If both of them failed to reproduce the bug, the bug was marked as reproducible. We spent $\approx$280 person-hours on the manual bug reproduction process.
@@ 4.5 Verification of Bug Reproduction
To verify the successful reproduction of bugs manifesting in error messages, we match the error messages with those found in the bug reports. On the other hand, to verify the reproduction of bugs with silent symptoms such as slow training or low accuracy, we adopt randomness in our edit actions such as input data generation, hyperparameter initialization, and neural network construction. For example, we generate random input data from the distributions or ranges suggested in the bug reports to avoid any unintended behaviours and to address the stochastic nature of these bugs. We also construct multiple neural networks of the same architecture but with different numbers of parameters or layers. Furthermore, we run the buggy code pertaining to silent bugs multiple times with different random values to ensure that the symptoms stem from the described bug rather than our added code. Following the successful reproduction of a bug, we capture various information related to each bug, such as the deep learning architecture involved, edit actions used to reproduce the bug, time taken to reproduce the bug, type of bug reproduced, and the type of information present in the bug report. All these data gathered during bug reproduction helped us identify key edit actions and information components to reproduce specific types of deep learning bugs (Step 4, Fig. 2).
@@ 4.6.4 Definitions for Apriori Algorithm
  • Transaction: A transaction is a record that contains one or more itemsets. In our study, we have two types of transactions as follows:
@@ 4.7 User Study
Instrument Design: We divided our questionnaire into three sections. We discuss them in detail below:
@@
Study Session: During our study session, each participant completes the following five tasks. First, the participant provides their demographic information. Second, the participant explains their daily challenges when reproducing deep learning bugs. Third, the participant reproduces two deep learning bugs and reports the edit actions and information they used to reproduce the bugs. Fourth, the participant also provides the rationale behind their used edit actions and component information to reproduce the bugs. Finally, the participant provides information about any other edit actions that they might have used to reproduce the bugs but are not covered in our study.
@@
Leveraging Associations to Produce Hints for User Study: In Step 3.6.6, we derived high-confidence associations to develop hints for the user study. These associations revealed edit actions and critical information needed to reproduce specific bugs. We then use this information to prepare the hints for the user study. In the user study, we do not directly provide the solution for reproducing the bugs. Instead, we give them a small hint (i.e., where to look in the bug report for one of the pieces of critical information or how to start the reproduction process), which helps them in bug reproduction. With this intervention, we plan to measure if our findings can help a person reproduce the bug more accurately or quickly.
@@ 5.1 RQ1: Which edit actions are crucial for reproducing deep learning bugs?
Obsolete Parameter Removal (A6) is a vital edit action to enhance the reproducibility of bugs where the differences in the library and framework versions may lead to problems. In this edit action, we remove all the obsolete parameters not supported by the newer version of the library or framework. Consider the issue reported in the Stack Overflow post (Issue #65992364). In this post, the reporter attempts to optimize an object detection model using ‘pytorch-mobile’, but the code snippet fails to optimize the model file size. The code snippet in the issue also contains the following line of code.
@@
However, according to the API documentation of PyTorch 1.6.0, the backend parameter was no longer supported. We removed the obsolete parameter and thus were able to reproduce the bug.
@@
tfhub_handle_preprocess = "https://tfhub.dev/tensorflow/bert_en_uncased_preprocess/3"
tfhub_handle_encoder = "https://tfhub.dev/tensorflow/small_bert/bert_en_uncased_L-4_H-512_A-8/1"
Version Migration (A10) is one of the commonly used edit actions for all types of bugs. In this edit action, we replace the outdated or deprecated versions of a library or framework with the latest version to reproduce the bug. Mondal et al. msr2019mondal introduced a similar edit operation known as ‘Code Migration’ in their work. This operation often confirms the validity of the bug by reproducing it with the latest version. While downgrading the library to an older version can help reproduce bugs relying on deprecated functionality, upgrading to a more recent version is still useful when downgrading is impossible due to dependency conflicts. Furthermore, upgrading to the recent version can also help determine if the bug persists across different versions, providing valuable information for developers. Moreover, when the developers set up the environment for the bug reproduction, installing the latest compatible versions of libraries and frameworks can be more straightforward than attempting to match the exact outdated versions mentioned in the bug report. In cases where downgrading is infeasible, upgrading to the recent version can help reproduce deep learning bugs. For example, the post on Stack Overflow (Issue #45711636) describes an issue with the CNN architecture constructed by the user. When the user passes the input through the CNN, they encounter a ValueError, which highlights a problem with the negative dimension value of the input. However, the code snippet provided in the issue uses Tensorflow 1.3.0. To reproduce this issue in TensorFlow 2.14.0, we carefully migrated the model built in TensorFlow 1.3.0 to the syntax of Tensorflow version 2.14.0. Specifically, we updated the Sequential, Conv2D and MaxPooling2D constructors to match the TensorFlow 2.14.0 API syntax. We also changed padding and pool size parameter schemes and removed the obsolete input shapes. Finally, we upgraded the model compilation and imported the required Tensorflow 2.14.0 modules. After modifying the code snippet, we could reproduce the bug successfully in our runtime environment.
@@ 5.3.1 Impact of Hints on Bug Reproducibility
To determine the impact of hints, we constructed a generalized linear model. This allowed us to test the statistical significance of multiple independent factors on bug reproducibility (i.e., dependent variable). The factors were experience with deep learning bug fixing, experience with deep learning, profession, and the presence/absence of hints.
@@
Variable Estimate Std. Error t value Pr $>$ ||||t||||
Intercept 0.4734 0.104 4.565 0.000
Hints 0.4011 0.086 4.651 0.000
Deep Learning Bug Fixing Experience 0.1085 0.142 0.763 0.445
Deep Learning Experience 0.0831 0.165 0.504 0.614
Profession (Academia/Industry) -0.2780 0.160 -1.737 0.082
Based on the regression results in Table 12, we can see that the presence of hints had a statistically significant positive effect on the reproducibility of deep learning bugs (p $<$ 0.001).
On the other hand, factors like experience with deep learning bug fixing, general deep learning experience, and profession (academia vs industry) did not significantly influence reproducibility. However, the participants’ profession (academia or industry) demonstrates a marginal negative effect on the reproducibility of deep learning bugs. (p = 0.082).
Overall, these results highlight that the presence of targeted hints has a significant positive impact on the reproducibility of deep learning bugs.
5.3.2 Impact of Developer Experience on Deep Learning Bug Reproducibility
To perform co-factor analysis, we construct another GLM model. We use deep-learning bug-fixing experience as a factor to evaluate its impact on the reproducibility of deep-learning bugs. The results of the analysis are shown in Table. 13.
Table 13: Simplified Model Coefficients
Variable Coef Std. Err z value P$>$||||z||||
Intercept 0.7839 0.300 2.610 0.009
DLBugFixExp_0[T.True] -0.2730 0.597 -0.457 0.648
DLBugFixExp_1[T.True] 0.7422 0.460 1.612 0.107
DLBugFixExp_2[T.True] 0.3147 0.651 0.484 0.629
The results from our generalised linear model suggest that years of experience in deep learning bug fixing do not have a statistically significant effect on the reproducibility of deep learning bugs. In particular, both developers with 1-5 years of experience and those with 5-10 years of experience achieved slightly higher bug reproducibility than those with less than 1 years of experience. However, neither of these experience levels have a statistically significant impact on bug reproducibility according to the model (p $>$ 0.05). On the other hand, as shown in Table. 12, the presence of targeted hints was found to have a statistically significant positive effect on bug reproducibility (p $<$ 0.001). Taken together, the above results indicate that the availability of specific guidance, rather than years of experience, is the major factor that can help the developers reproduce deep-learning bugs.

### v3

@@ 4.1 Selection of Data Sources
We select Stack Overflow as a primary data source for our study. It is the largest programming Q&A site for programming topics containing over 24 million questions and 35 million answers SOUsageData . Thus, Stack Overflow could be a potential source for deep learning (DL) bugs. Developers often submit their encountered problems on Stack Overflow, when building their deep learning applications. According to a recent work zhao2021state , Stack Overflow contains at least 30 topics and 80K posts related to deep learning issues, which makes it a potential source for our data. We also select the benchmark dataset of Moravati et al. defects4ml (Defects4ML in Fig. 2) containing 100 DL bugs as our second source of data. Specifically, we use 63 GitHub issues and 37 Stack Overflow posts from the Defects4ML dataset in our study.
@@
We combine all four filtration criteria above, construct a SQL query, and then execute the query against the Stack Exchange Data Explorer StackExchange . The SQL query is available in our replication package replicationPackage . After this filtration step, we obtained a total of 279 posts for Keras, 1,433 posts for Tensorflow, and 1,700 posts for PyTorch. To ensure the validity of our chosen filtration criteria and confirm that the posts are related to DL bugs, we conducted a follow-up manual analysis on a representative sample of the filtered posts. We determined the appropriate sample size using Cochran’s sample size formula cochran1977sampling :
$n=\frac{Z^{2}\cdot p\cdot(1-p)}{e^{2}}$
Where:
  • $n$ = sample size
  • $Z$ = Z-value for the desired confidence level (1.645 for 90% confidence)
  • $p$ = population proportion (assumed to be 0.5 for maximum variability)
  • $e$ = margin of error (0.1 or 10%)
After plugging in the values:
$n=\frac{1.645^{2}\cdot 0.5\cdot(1-0.5)}{0.1^{2}}=67.65$
Therefore, a representative sample of 67 posts was needed to achieve 90% confidence with a 10% margin of error. We chose these parameters to balance the precision of the estimates with the manual effort required for the analysis. Moreover, higher confidence levels or lower margins of error would have necessitated significantly larger sample sizes and analysis time as follows:
  • 95% confidence, 5% margin of error: $n=346$ posts (86.5 hours)
  • 99% confidence, 1% margin of error: $n=2,832$ posts (708 hours)
For each of the 67 sampled posts, we analyzed the entire Q&A discussion thread from Stack Overflow and the corresponding code changes in GitHub. This manual analysis took approximately 15 minutes per post, resulting in a total of around 17 hours of effort. We found that $\approx$96of the sampled posts discussed deep learning bugs, confirming the validity of our filtration criteria (Steps 1-2, Fig. 2).
@@ 4.4 Qualitative Analysis
Before conducting any qualitative analysis, we first determine if each post targets a deep learning bug. If a post is not related to a deep learning bug, we exclude it from further analysis. We also excluded the posts not reporting the evaluation metrics of the buggy model from our study. Without these metrics, it was impossible to determine if a bug was successfully reproduced or not. This two-step filtering process helped our analysis focus on relevant and measurable deep-learning bug reports.
Once we confirm that the post is relevant to a DL bug, we check if each selected post accurately represents the type of bug it is assigned to. To validate the bug categorization, we manually analyze each selected Stack Overflow post and review the issue description, stack trace, observed behaviour, expected behaviour, and accepted answer. This ensures that the tag-based selection of posts from Stack Overflow does not affect our results’ validity. We perform this manual verification since the Stack Overflow posts might lack an independently verified category label. In contrast, the GitHub issues in the benchmark dataset by Morovati et al. defects4ml already contain a validated category label.
4.4.1 Reproducing Deep Learning Bugs from Stack Overflow
Once we confirm the category of each post or issue, we follow a two-step approach to reproduce the deep-learning bugs from Stack Overflow posts. First, we gather complementary information about each bug, such as the dataset, code snippet, library versions, and the framework used. Second, we attempt to reproduce the bug using the code snippet and supporting data (e.g., dataset information, environment configurations, hyperparameters, and training logs).
4.4.2 Reproducing Bugs from Github Issues
To reproduce the bugs from GitHub issues, we employed a systematic approach. First, we located the bug-inducing commit of a bug using the commit ID provided by the benchmark. Next, we cloned the corresponding repository and checked out the buggy version of the code using the commit ID. We then set up the development environment according to the reported bug, which may include installing the required version of the programming language and necessary dependencies or libraries, configuring environment variables and setting up any necessary databases, services, or external dependencies the code requires. If applicable, we applied necessary edit actions to the code, such as modifying specific lines or adding/removing code snippets, to trigger the buggy behaviour. We then ran the updated code snippet or the specific part of the codebase where the bug was expected to occur. To verify the presence of a bug, we compared the observed behaviour with the reported buggy behaviour and ran the test cases from the existing benchmark (Defects4ML). Finally, if the bug was reproduced, we recorded the edit actions used and critical information necessary for the bug reproduction.
4.4.3 Agreement Analysis
The first author and one independent collaborator conducted the bug reproduction process. We first reproduced 10 bugs from our dataset and achieved a Cohen Kappa of 54.5%. Then, we had two meetings to identify the main reasons for our disagreements and resolved them. In the next round, we reproduced 10 more bugs and achieved a Cohen Kappa of 89.1%, which is considered an almost perfect agreement mchugh2012interrater . After achieving an almost perfect agreement, the first author reproduced the remaining bugs (i.e., 128) while the independent collaborator checked the reproduction, achieving an average Cohen Kappa of 85.4%. If the first author fails to reproduce the bug within 60 minutes, the independent collaborator also attempts to reproduce the bug. If both of them failed to reproduce the bug, the bug was marked as irreproducible. We spent $\approx$280 person-hours on the manual bug reproduction process.
@@ 4.5 Verification of Bug Reproduction
To verify the successful reproduction of bugs, we employed different strategies depending on the nature of the bugs.
4.5.1 Explicit Bugs
An explicit bug results in an error message or exception. To verify the reproduction of this bug, we adopted a straightforward approach. We extracted the error message from the bug report and attempted to reproduce the bug under the reported conditions. We considered the bug reproduction successful when the observed error message matched the error message in the bug report.
4.5.2 Silent Bugs
Silent bugs are also referred to as functional or numerical errors. They do not result in system crashes or hangs and do not display error messages, but they lead to incorrect behaviour tambon2024silent . We adopted a comprehensive approach to verify the reproduction of silent bugs, as they often manifest subtly through silent issues like slow training or low accuracy.
We reproduced the silent bugs from Stack Overflow posts and GitHub issues. Subsequently, we used the evaluation metrics reported in the original posts or issues as the ground truth to verify the buggy behaviour. In cases where the code snippet was incomplete, we applied our edit actions to make it compilable, executable, and runnable.
To determine if a bug was successfully reproduced, we followed these steps:
  • We executed the modified code snippet five times, each time with a different random seed, and calculated the average evaluation metric across these runs.
  • We compared the average evaluation metric to the reported evaluation metric of the buggy model.
  • If the average evaluation metric was within a 5% error margin of the reported metric, we considered the bug to be reproduced.
We selected the 5% threshold for error margin based on the existing literature pham2020problems ; alahmari2020challenges . Pham et al.pham2020problems found that implementation-level non-determinism could account for $\approx$3variance in the training and evaluation metrics of DL models, often caused by factors such as parallel processing issues, automatic selection of primitive operations, task scheduling, and floating-point precision differences. Furthermore, Alahmari et al.alahmari2020challenges demonstrated that the variance in evaluation metrics could vary from 3% to 7% for models trained using the same dataset and code. Thus, our 5% threshold accounts for the inherent variability in deep-learning models while still maintaining a reasonable standard for bug reproduction. This systematic approach allows us to verify the successful reproduction of silent bugs, ensuring the reliability of our findings.
4.5.3 Information Collection During Verification
Following the successful reproduction of any bug above, we capture various information related to each bug, such as the deep learning architecture involved, edit actions used to reproduce the bug, time taken to reproduce the bug, type of bug reproduced, and the type of information present in the bug report. All these data gathered during bug reproduction helped us identify key edit actions and information components to reproduce specific types of deep learning bugs (Step 4, Fig. 2).
@@ 4.6.4 Definitions for Apriori Algorithm
  • Transaction: A transaction is a record that contains one or more items. In our study, we have two types of transactions as follows:
@@ 4.7 User Study
Instrument Design: We used Opinio, an online survey tool recommended by our institution, to construct and distribute our questionnaire. Opinio enabled us to track the time spent by the participants on each individual question, which proved to be useful for our further analysis. The use of Opinio also did not require any additional effort from the participants, which made it a suitable choice for our user study. We divided our questionnaire into three sections. We discuss them in detail below:
@@
Study Session: During our study session, each participant completes the following five tasks. First, the participant provides their demographic information. Second, the participant explains their daily challenges when reproducing deep learning bugs. Third, each participant reproduces two deep learning bugs and self-reports the edit actions and information they used to reproduce the bugs. Fourth, the participant also provides the rationale behind their self-reported edit actions and component information used to reproduce the bugs. Finally, the participant provides information about any other edit actions that they might have used to reproduce the bugs but are not covered in our study.
@@
Leveraging Associations to Produce Hints for User Study: We use high-confidence associations from our RQ2 (Step 3.6.6) to develop hints for our user study. These associations revealed the edit actions and critical information needed to reproduce specific types of bugs. We use the steps below to construct the hints in a systematic way for our user study.
  1. 1.
    Bug Categorization:
    1. (a)
      Analyze the information found in a bug report to identify the specific type or category of the bug.
    2. (b)
      Use the identified bug category to retrieve relevant component information and edit actions.
  2. 2.
    Retrieve component information and edit actions:
    1. (a)
      Based on the identified bug category, retrieve the top 3 components and top 5 edit actions from the findings of RQ2.
    2. (b)
      The component information includes the most important aspects required to reproduce or understand a bug (e.g., shape of input data, training code, error messages).
    3. (c)
      The edit actions refer to the most frequently associated actions to a specific bug category.
  3. 3.
    Determining the Most Relevant Statement: We also collect the most relevant statement from the bug description as follows.
    1. (a)
      Split the bug report’s text into a list of sentences using ’.’ as the delimiter.
    2. (b)
      Collect a list of key phrases for the components retrieved in Step 2(a). These key phrases are derived from our qualitative analysis of the bug reports and can be found in the replication package replicationPackage .
    3. (c)
      Generate the Sentence-BERT embeddings for every statement in the bug report and the keywords above using the ’sentence-transformers/all-MiniLM-L6-v2’ pre-trained model.
    4. (d)
      For each statement in the bug report, calculate the cosine similarity between the statement embedding and the embeddings of the keywords.
    5. (e)
      Identify the relevant statement for the most prevalent component (i.e., top component from Step 2(a)) based on their cosine similarity score.
    6. (f)
      Prepare a context-specific hint that guides the user to a particular statement using the following template: ‘Focus on the statement: $<$Statement extracted in Step 3(e)$>$
  4. 4.
    Hint Formulation:
    1. (a)
      Combine the retrieved component information and edit actions from the research findings with the context-specific hint to formulate the complete hint for the user study.
    2. (b)
      Use the template below to formulate the hint for each bug.
Template for Hint Generation Hints 1. $<$CI1$>$, $<$CI2$>$, $<$CI3$>$ can be useful information for reproducing the bug. 2. $<$EA1$>$, $<$EA2$>$, $<$EA3$>$, $<$EA4$>$, $<$EA5$>$ can be useful edit actions for reproducing the bug. 3. Focus on the Statement: $<$Most Relevant Statement from the SO Post$>$
We ensure a systematic formulation of hints for our study by following the steps above. This approach incorporates high-confidence associations from Step 3.6.6 and the bug description for individual bug. With this intervention, we plan to measure if and how our recommended information help participants reproduce the bugs more accurately or quickly. The code and results for the hint formulation are available in our replication package replicationPackage .
@@ 5.1 RQ1: Which edit actions are crucial for reproducing deep learning bugs?
Obsolete Parameter Removal (A6) is a crucial edit action that enhances bug reproducibility when differences in library and framework versions cause compatibility issues. It involves removing obsolete parameters no longer supported by newer versions. Frequent updates to deep learning frameworks and APIs can cause breaking changes, i.e., the code written for older versions of the framework is incompatible with newer versions. Developers reproducing these bugs often face challenges when the bug report’s environment is significantly older than their current working environment, and downgrading to match those outdated versions is not always feasible due to various constraints. In some cases, the bug report uses significantly outdated versions that are no longer supported or maintained by the framework developers. Additionally, the developer’s current project may have dependencies that require newer versions of Python or the frameworks, making it impractical to downgrade. Furthermore, organizations or development teams may have policies in place that mandate using the latest stable versions for security, performance, and maintainability reasons, preventing the use of older versions.
In such cases, developers must port the bug-reproducing code to their current environment, where Obsolete Parameter Removal edit action proves highly beneficial. It makes the code compatible with newer framework versions while preserving the bug-reproducing behaviour. Consider the issue reported in the Stack Overflow post (Issue #65992364). In this post, the reporter attempts to optimize an object detection model using ‘pytorch-mobile’, but the code snippet fails to optimize the model file size. The code snippet in the issue also contains the following line of code.
@@
However, according to the API documentation of PyTorch 1.6.0 pytorch1.6 , the backend parameter was no longer supported. We removed the obsolete parameter and thus were able to reproduce the bug.
@@
tfhub_handle_preprocess = ‘‘https://tfhub.dev/tensorflow/bert_en_uncased_preprocess/3"
tfhub_handle_encoder = ‘‘https://tfhub.dev/tensorflow/small_bert/bert_en_uncased_L-4_H-512_A-8/1"
Version Migration (A10) is a vital edit action for reproducing bugs in deep learning frameworks and libraries. It involves adapting code written for older versions to the latest version, confirming the same bug’s presence in the current environment. Mondal et al. msr2019mondal introduced a similar concept called ‘Code Migration’. Similar to Obsolete Parameter Removal, Version Migration aims to address compatibility issues caused by the updates in deep learning frameworks and APIs. While Obsolete Parameter Removal focuses on removing deprecated parameters, Version Migration contains additional modifications required to make the code follow the newer version’s syntax, APIs, and functionalities. This may involve updating function calls, class constructors, import statements, and other code elements affected by the version changes.
The necessity for Version Migration often arises when the versions used in the bug report are significantly outdated compared to the developer’s current working environment. Bug reports may contain code snippets written for older framework versions that are no longer actively maintained or supported by the developers. In such cases, downgrading to the exact reported version is not feasible, as it would require reverting to unsupported and potentially insecure versions. Additionally, the developer’s current project may have dependencies that require newer versions of Python or deep learning frameworks, making it impractical to downgrade. Furthermore, organizations or development teams may have policies in place that mandate using the latest stable versions for security, performance, and maintainability reasons, preventing the use of older versions. To demonstrate the utility of the Version Migration, let us consider the issue reported in the post on Stack Overflow (Issue #45711636). The post describes an issue with the CNN architecture constructed by the user. When the user passes the input through the CNN, they encounter a ValueError, which highlights a problem with the negative dimension value of the input. However, the code snippet provided in the issue uses Tensorflow 1.3.0. To reproduce this issue in TensorFlow 2.14.0, we carefully migrated the model built in TensorFlow 1.3.0 to the syntax of Tensorflow version 2.14.0. Specifically, we updated the Sequential, Conv2D and MaxPooling2D constructors to match the TensorFlow 2.14.0 API syntax. We also changed padding and pool size parameter schemes and removed the obsolete input shapes. Finally, we upgraded the model compilation and imported the required Tensorflow 2.14.0 modules. After modifying the code snippet, we successfully reproduced the bug in our runtime environment. We verified the reproduction by comparing the error message from our updated code snippet with the one reported in the Stack Overflow post.
@@ 5.3.1 Impact of Hints on Bug Reproducibility
To determine the impact of our hints on bug reproducibility, we constructed a generalized linear model (GLM) using the Binomial family with a logit link function nelder1972generalized . This allowed us to test the statistical significance of multiple independent factors on bug reproducibility (i.e, our dependent variable). The factors were experience with deep learning bug fixing, experience with deep learning, profession, and the presence/absence of hints, with ‘Hints’ being our main factor of interest. We chose Odds Ratio as our effect size metric for two main reasons. First, it has been commonly used in similar studies within software engineering research, as demonstrated by Ceccato et al. ceccato2014family . Second, the Odds Ratio is appropriate for a logistic regression-based model with a binary target variable, which we used in our study.
@@
Variable Estimate Std. Error z value Pr $>$ ||||z|||| Effect Size (OR)
Intercept -3.4229 1.898 -1.803 0.071 -
DLBugFixExp_0 0.2251 0.856 0.263 0.793 1.252493
DLBugFixExp_1 -0.2449 0.564 -0.435 0.664 0.782786
DLBugFixExp_2 -3.4032 2.283 -1.491 0.136 0.033268
Hints 3.0899 0.991 3.118 0.002 21.974308
DLExp 2.5448 1.725 1.475 0.140 12.740844
Field 0.1915 1.145 0.167 0.867 1.211112
Based on the regression results in Table 12, we can see that the presence of hints had a statistically significant positive effect on the reproducibility of deep learning bugs (p = 0.002 $<$ .05). The effect size, as measured by the odds ratio, indicates that the presence of hints increases the odds of reproducing a deep learning bug by a factor of 21.97 compared to the absence of hints. The intercept term, with an estimate of -3.4229 (p = 0.071), represents the baseline probability of reproducing a deep learning bug when no hints are provided, the participant has no experience with deep learning bug fixing, deep learning in general, and their profession is not considered. Moreover, as shown in Table. 12, factors like experience with deep learning bug fixing, general deep learning experience, and profession (academia vs industry) did not significantly influence reproducibility.
Overall, the above results suggest that the presence of targeted hints positively impacts the reproducibility of deep learning bugs with a statistically significant margin and a large effect size. Other factors, such as experience and profession, do not play a significant role in bug reproducibility.